Find the first five terms of the sequence for which A1=5 and An+1=3An+1
I'm not sure if I started this right. 3(5)+1=16
Is this right? Could you please explain how to do this problem.
Thanks

Answer by oscargut(2103)

A1=5
A2= 3(A1)+1=3(5)+1=16 (you were right)
A3= 3(A2)+1=3(16)+1=49
A4=3(A3)+1=3(49)+1=148
A5=3(A4)+1=3(148)+1=445